OpenStudy (anonymous):

The measure of two complementary angles are in the ratio 1:4. What are the degree measures of the two angles

OpenStudy (raden):

complementary means add up both angles to 90 degrees let the first angle is a = x and the other is b = 4x (see that a : b = 1 : 4) now take an equation : a + b = 90 x + 4x = 90 5x = 90 x = 90/5 x = 18 so, a = x = 18 and b = 4x = 4(18) = 72